OpenStudy (anonymous):

What number should be added to -85 to get a sum of -2?

OpenStudy (ash2326):

Sum =-85+x -2=-85+x 85-2=x or x=83

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

when numbers are positive to get to negative subtract ( - ) ______________________ / V <------|---------|---------|-------> - 5 0 5 ^ L___________________> ( + ) when numbers are negative to go to the positive side add